A small, square loop carries a 38 A current. The on-axis magnetic field strength 44 cm from the loop is 4.7 nT .

What is the edge length of the square?

Express your answer to two significant figures and include the appropriate units.
Answer #1

47T μ,21 AAe 4.7 × 10-9) 4π(0.44 (47x10) 4(0.44) (4π×10-7) 2 (38) = 5.3×10-5 m2 side=VA=7.2x10-3 m-7.2 mm